# supplement

### v1

@@ Appendix A Additional related work
The theoretical foundations of conformal prediction were established by [31, 30, 44, 57]. This framework has since been extended to accommodate covariate shift by applying weights to the distribution of conformity scores [54]. Similar weighting strategies have also been employed to integrate active learning into conformal prediction [45]. Building further on these weighted approaches, conformal prediction methods have recently been adapted for survival analysis tasks [5, 15, 9].
@@
Several recent works adapt conformal ideas to off-the-shelf LLMs. For instance, [60] adapts conformal prediction to open-ended generation by integrating a self–consistency–based uncertainty measure, achieving strict correctness coverage across seven popular LLMs and diverse domains. The work of [58] further extends this by explicitly inserting the ground truth into candidate outputs and using logit-based nonconformity scores to guarantee coverage over a wider range of error rates, though its reliance on logits raises calibration concerns when only API-level access is available. Additionally, Conformal Language Modeling has been developed in [46] to calibrate both stopping and rejection rules to guarantee coverage in open-domain QA and summarization. The authors of [36] distinguish uncertainty from confidence using semantic dispersion metrics, enabling selective generation of low-confidence outputs. To refine validity under practical constraints, [7] extended conditional conformal methods to incorporate utility-based guarantees and differentiable filtering. [41] introduced “conformal factuality,” constructing entailment-based uncertainty sets that deliver high-probability correctness. The approach proposed in [16] provides a guarantee for aligned responses in QA and radiology applications.
@@
Active learning methods aim to select the most informative samples for labeling, improving model accuracy under tight budgets [52]. This was formalized in [67] for statistical inference tasks, where one prioritizes labeling points with high model uncertainty while maintaining valid confidence intervals. Our Optimized method relates to this idea as it derives sample probabilities that directly minimize the variance of the augmented inverse-propensity-weighting estimator.
@@
Optimal experimental design allocates observations to minimize estimator variance under a given model. In [17], the authors develop a two-stage adaptive design that chooses propensity scores to minimize the asymptotic variance bound in treatment-effect estimation, and the approach presented in [37] presents simple rules of thumb for efficient designs under resource constraints. Analogously, our Optimized method treats the sampling rule $\pi(x)$ as a design variable and solves for the allocation that minimizes the miscoverage estimation variance.
@@
The survival analysis problem was first formalized in biostatistics and reliability engineering. The work in [27] introduces a nonparametric estimator of the survival function for right-censored data, laying the groundwork for all subsequent methods. The proportional hazards model was later developed, formalizing a semiparametric regression framework that relates covariates to an event’s hazard rate without specifying its baseline form. As an alternative formulation, accelerated failure time (AFT) [35] models assume a parametric form for the log-survival time, allowing direct modeling of time-to-event via common distributions such as Weibull or log-normal. More recently, machine-learning approaches have been developed for survival prediction. Random survival forests leverage ensemble tree methods to nonparametrically estimate cumulative hazard functions [22]. DeepSurv employs a neural-network approximation of the Cox partial likelihood to capture complex covariate interactions [28], and Deep Cox Mixtures generalizes this by modeling mixtures of proportional hazards to flexibly adapt to heterogeneous subpopulations [42].
@@ Appendix B Calibration algorithms
B.1 Naive calibration
Algorithm 1 in this section outlines the Naive method to construct reliable LPBs.
Algorithm 1 Naive Calibration
0:  Calibration data $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, generative model $\mathcal{G}(\cdot)$, audit function $\texttt{Audit}(\cdot)$, pre-trained quantile regression model $\{\hat{q}_{\tau}(\cdot)\}_{\tau\in\mathcal{T}}$, target miscoverage rate ${\alpha}$, and total budget $B$.
1:  Sample $C_{i}\sim\mathrm{Geom}(n/B)$ for all ${i\in{\mathcal{I}_{2}}}$
2:  for each $i\in{\mathcal{I}_{2}}$ do
3:     Initialize $j\leftarrow 0$
4:     repeat
5:        $j\leftarrow j+1$
6:        Generate and evaluate the output’s safety: $Y_{i}^{j}\leftarrow\texttt{Audit}(X_{i},\mathcal{G}(X_{i}))$
7:     until $Y_{i}^{j}=1$ or $C_{i}=j$
8:     Set $\tilde{T}_{i}\leftarrow j$
9:  end for
10:  $w_{\tau}(X_{i})\leftarrow\frac{1}{\mathbb{P}[\hat{q}_{\tau}(X_{i})\leq C_{i}|X_{i}]},\ i\in{\mathcal{I}_{2}}$ // compute the weights for the geometric $C_{i}$
11:  for $\tau\in\mathcal{T}$ do
12:     $\hat{\alpha}(\tau)\leftarrow\displaystyle\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w_{\tau}(X_{i})\;\mathbb{I}\bigl{\{}\hat{q}_{\tau}(X_{i})\leq C_{i}\bigr{\}}\;\mathbb{I}\bigl{\{}T_{i}<\hat{q}_{\tau}(X_{i})\bigr{\}}$ // miscoverage rate estimator
13:  end for
14:  $\hat{\tau}\leftarrow\sup\Bigl{\{}\tau\in\mathcal{T}:\sup_{\begin{subarray}{c}\tau^{\prime}\in\mathcal{T}\\ \tau^{\prime}\leq\tau\end{subarray}}\hat{\alpha}(\tau^{\prime})\leq{\alpha}\Bigr{\}}$   // calibrated quantile level
15:  return  Lower predictive bound (LPB) for a test point $X_{\textup{test}}=x$, given by $\hat{L}(x)=\hat{q}_{\hat{\tau}}(x)$
B.2 Adaptive budget calibration
@@
where $\lambda>0$ is the Lagrange multiplier. On the open domain $(0,1]^{n}$, the objective is strictly convex and differentiable, so the single stationary point satisfying the first-order conditions is in fact the unique global minimizer. Taking partial derivatives,
@@
Now, define the budget-usage function
@@
Algorithm 2 Prompt-Adaptive Budget Calibration: Adaptive method
0:  Calibration data $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, generative model $\mathcal{G}(\cdot)$, audit function $\texttt{Audit}(\cdot)$, pre-trained quantile regression model $\{\hat{q}_{\tau}(\cdot)\}_{\tau\in\mathcal{T}}$, target miscoverage rate ${\alpha}$, prior quantile ${\tau_{\textup{prior}}}$, calibration mode (Basic, Trimmed, Optimized), quantile trimming threshold $M$, total budget $B$.
1:  if calibration mode is Basic then
2:     $\{\hat{f}_{\tau}(X_{i})\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}\leftarrow\{\hat{q}_{\tau}(X_{i})\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}$
3:     $\{\pi_{i}\}\leftarrow\{\min(\tfrac{B}{|{\mathcal{I}_{2}}|\;\hat{f}_{\tau_{\textup{prior}}}(X_{i})},1)\}$   // per-prompt evaluation probability (Section 5.1)
4:  else if calibration mode is Trimmed then
5:     $\{\hat{f}_{\tau}(X_{i})\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}\leftarrow\{\min{(\hat{q}_{\tau}(X_{i}),M)}\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}$   // trim the quantile est. (Section 5.2)
6:     $\{\pi_{i}\}\leftarrow\{\min(\tfrac{B}{|{\mathcal{I}_{2}}|\;\hat{f}_{\tau_{\textup{prior}}}(X_{i})},1)\}$   // per-prompt evaluation probability (Section 5.1)
7:  else if calibration mode is Optimized then
8:     $\{\hat{f}_{\tau}(X_{i})\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}\leftarrow\{\min{(\hat{q}_{\tau}(X_{i}),M)}\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}$   // trim the quantile est. (Section 5.2)
9:     $\{\pi_{i}\}\leftarrow$ Algorithm 4 applied with $\{\hat{f}_{\tau_{\textup{prior}}}(X_{i})\}_{i\in{\mathcal{I}_{2}}}$   // optimized per-prompt evaluation                       probability (Section 5.3)
10:  end if
11:  $\{(\tilde{T}_{i},C_{i})\}_{i\in{\mathcal{I}_{2}}}\leftarrow$ Algorithm 3 applied to $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$ with $\{\pi_{i}\}_{i\in{\mathcal{I}_{2}}}$ and $\{\hat{f}_{\tau_{\textup{prior}}}(X_{i})\}_{i\in{\mathcal{I}_{2}}}$
12:  $w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\leftarrow\frac{1}{\pi_{i}},\ i\in{\mathcal{I}_{2}}$
13:  for $\tau\in\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$ do
14:     $\hat{\alpha}(\tau)\leftarrow\displaystyle\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\;\mathbb{I}\bigl{\{}\hat{f}_{\tau}(X_{i})\leq C_{i}\bigr{\}}\;\mathbb{I}\bigl{\{}T_{i}<\hat{f}_{\tau}(X_{i})\bigr{\}}$ // miscoverage est.
15:  end for
16:  $\hat{\tau}\leftarrow\sup\Bigl{\{}\tau\in\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]:\sup_{\begin{subarray}{c}\tau^{\prime}\in\mathcal{T}\\ \tau^{\prime}\leq\tau\end{subarray}}\hat{\alpha}(\tau^{\prime})\leq{\alpha}\Bigr{\}}$   // calibrated quantile level
17:  return  Lower predictive bound (LPB) for a test point $X_{\textup{test}}=x$, given by $\hat{L}(x)=\hat{f}_{\hat{\tau}}(x)$
@@
0:  Calibration data $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, generative model $\mathcal{G}(\cdot)$, max per-prompt budget $\{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\}_{i\in\mathcal{I}}$, per-prompt evaluation probability $\{\pi_{i}\}_{i\in\mathcal{I}}$, audit function $\texttt{Audit}(\cdot)$
@@
4:     Initialize $j\leftarrow 0$
5:     repeat
6:        $j\leftarrow j+1$
7:        Generate and evaluate the output’s safety: $Y_{i}^{j}\leftarrow\texttt{Audit}(X_{i},\mathcal{G}(X_{i}))$
8:     until $Y_{i}^{j}=1$ or $C_{i}=j$
9:     Set $\tilde{T}_{i}\leftarrow j$
10:  end for
11:  return  $\{(\tilde{T}_{i},C_{i})\}_{i\in\mathcal{I}_{2}}$
@@
7:     $\lambda^{*}\leftarrow$ bisection on $[\lambda_{\textup{low}},\lambda_{\textup{high}}]$ to solve $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}^{*}(\lambda)=B$ within $\epsilon$
@@ Appendix C A deeper look into $\hat{\alpha}(\tau)$
In Section 4.1, we introduced the miscoverage estimator $\hat{\alpha}(\tau)$. In this section, we explore its statistical properties in detail. We begin by deriving its expectation and variance:
@@
See Appendix D.1 for the proof. We remark that the conditional independence and pair-wise independence assumptions in Proposition C.1 hold for all variants of our method from Sections 4 and 5. This proposition shows that $\hat{\alpha}(\tau)$ is an unbiased miscoverage estimator, resulting in a useful approximation for the miscoverage estimation variance. Under a simplified assumption that, for a fixed $\tau$, each calibration point $i\in{\mathcal{I}_{2}}$ is miscovered at the same constant rate conditional on $\{X_{j}\}_{j\in\mathcal{I}_{2}}$, we have the following result.
@@ Appendix D Proofs
Observe that if $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\leq B$, then $\pi_{i}^{*}=1$ for all $i\in{\mathcal{I}_{2}}$ is the optimal solution which satisfies the constraint. We now turn to consider the case where $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$. We first show that there exists a unique solution $\lambda^{*}$. Observe that $U:(0,\infty)\to(0,\sum_{i}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})]$ is continuous and strictly decreasing, with $\lim_{\lambda\to 0^{+}}U(\lambda)=\sum_{i}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$ and $\lim_{\lambda\to\infty}U(\lambda)=0$. By the intermediate-value theorem there is a unique $\lambda^{*}>0$ such that $U(\lambda^{*})=B$. Next, the output of Algorithm 4 is indeed the solution $\lambda^{*}$ since $U$ is a monotone function, and thus the bisection approach is valid in this case. Therefore, $\pi^{*}(\lambda)$ is the correct optimal solution. ∎
@@
Inductive step. We now suppose that the claim holds for any weight vector with a sum $B-1<S\leq M|{\mathcal{I}_{2}}|$, and show that it holds for any vector with a sum $S-1$. Given a weight vector $\{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\}_{i\in{\mathcal{I}_{2}}}$ with a sum $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})=S-1$, we compose a new vector with a sum $S$, as follows. In $\{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\}_{i\in{\mathcal{I}_{2}}}$, there exists an index $i_{0}$ such that $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})\leq M-1$. The new vector with a sum $S$ is defined by:
@@
Since $U^{\prime}$ is continuous and decreasing on $(0,\infty)$ with $\lim_{\mu\to 0^{+}}U^{\prime}(\mu)=\sum_{i}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$ and $\lim_{\mu\to\infty}U^{\prime}(\mu)=0$, the intermediate-value theorem guarantees a unique $\lambda^{\prime}$ with $U^{\prime}(\lambda^{\prime})=B$. Moreover, since $U^{\prime}$ is decreasing, and $U^{\prime}(\lambda)\geq B=U^{\prime}(\lambda^{\prime})$ we get $\lambda^{\prime}\geq\lambda$. Define $\pi_{i}$ the optimal solution of (8) with $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$ and $\pi^{\prime}_{i}$ the optimal solution with $\hat{f}^{\prime}_{{\tau_{\textup{prior}}}}(X_{i})$.

### v4

@@ Appendix A Additional related work
The theoretical foundations of conformal prediction were established by Kolmogorov (1983; 1968), Papadopoulos et al. (2002), Vovk et al. (2005). This framework has since been extended to accommodate covariate shift by applying weights to the distribution of conformity scores (Tibshirani et al., 2019). Similar weighting strategies have also been employed to integrate active learning into conformal prediction (Prinster et al., 2024). Building further on these weighted approaches, conformal prediction methods have recently been adapted for survival analysis tasks (Candès et al., 2023, Gui et al., 2024, Davidov et al., 2025).
@@
Several recent works adapt conformal ideas to off-the-shelf LLMs. For instance, Wang et al. (2024) adapts conformal prediction to open‐ended generation by integrating a self–consistency–based uncertainty measure, achieving strict correctness coverage across seven popular LLMs and diverse domains. The work of Wang et al. (2025) further extends this by explicitly inserting the ground truth into candidate outputs and using logit‐based nonconformity scores to guarantee coverage over a wider range of error rates, though its reliance on logits raises calibration concerns when only API‐level access is available. Additionally, Conformal Language Modeling has been developed by Quach et al. (2024) to calibrate both stopping and rejection rules to guarantee coverage in open-domain QA and summarization. Lin et al. (2024) distinguish uncertainty from confidence using semantic dispersion metrics, enabling selective generation of low-confidence outputs. To refine validity under practical constraints, Cherian et al. (2024) extended conditional conformal methods to incorporate utility-based guarantees and differentiable filtering. Mohri & Hashimoto (2024) introduced “conformal factuality,” constructing entailment-based uncertainty sets that deliver high-probability correctness. The approach proposed in (Gui et al., 2024) provides a guarantee for aligned responses in QA and radiology applications.
@@
Active learning methods aim to select the most informative samples for labeling, improving model accuracy under tight budgets (Settles, 2010). This was formalized in (Zrnic & Candès, 2024) for statistical inference tasks, where one prioritizes labeling points with high model uncertainty while maintaining valid confidence intervals. Our Optimized method relates to this idea as it derives sample probabilities that directly minimize the variance of the augmented inverse-propensity-weighting estimator.
@@
Optimal experimental design allocates observations to minimize estimator variance under a given model. In (Hahn et al., 2011), the authors develop a two-stage adaptive design that chooses propensity scores to minimize the asymptotic variance bound in treatment-effect estimation, and the approach presented in (List et al., 2011) presents simple rules of thumb for efficient designs under resource constraints. Analogously, our Optimized method treats the sampling rule $\pi(x)$ as a design variable and solves for the allocation that minimizes the miscoverage estimation variance.
@@
The survival analysis problem was first formalized in biostatistics and reliability engineering. The work in (Kaplan & Meier, 1958) introduces a nonparametric estimator of the survival function for right‐censored data, laying the groundwork for all subsequent methods. The proportional hazards model was later developed, formalizing a semiparametric regression framework that relates covariates to an event’s hazard rate without specifying its baseline form. As an alternative formulation, accelerated failure time (AFT) (Lin et al., 1998) models assume a parametric form for the log‐survival time, allowing direct modeling of time-to-event via common distributions such as Weibull or log-normal. More recently, machine-learning approaches have been developed for survival prediction. Random survival forests leverage ensemble tree methods to nonparametrically estimate cumulative hazard functions (Ishwaran et al., 2008). DeepSurv employs a neural-network approximation of the Cox partial likelihood to capture complex covariate interactions (Katzman et al., 2018), and Deep Cox Mixtures generalizes this by modeling mixtures of proportional hazards to flexibly adapt to heterogeneous subpopulations (Nagpal et al., 2021).
@@ Appendix B Calibration algorithms
where $\lambda>0$ is the Lagrange multiplier. On the open domain $(0,1]^{n}$, the objective is strictly convex and differentiable, so the single stationary point satisfying the first‐order conditions is in fact the unique global minimizer. Taking partial derivatives,
@@
Now, define the budget‐usage function
@@
Algorithm 1 Generate Calibration Responses
0: Calibration data $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, Generative model $\mathcal{G}(\cdot)$, audit function $\texttt{Audit}(\cdot)$, Censoring times $\{C_{i}\}_{i\in{\mathcal{I}_{2}}}$.
1:for each $i\in\mathcal{I}$ do
2:  Initialize $j\leftarrow 0$
3:  repeat
4:   $j\leftarrow j+1$
5:   Generate and evaluate the output’s safety: $Y_{i}^{j}\leftarrow\texttt{Audit}(X_{i},\mathcal{G}(X_{i}))$
6:  until $Y_{i}^{j}=1$ or $C_{i}=j$
7:  Set $\tilde{T}_{i}\leftarrow j$
8:end for
9:return $\{\tilde{T}_{i}\}_{i\in\mathcal{I}_{2}}$
@@
0: Calibration data $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, generative model $\mathcal{G}(\cdot)$, max per-prompt budget $\{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\}_{i\in\mathcal{I}}$, per-prompt evaluation probability $\{\pi_{i}\}_{i\in\mathcal{I}}$, audit function $\texttt{Audit}(\cdot)$.
@@
4:end for
5: Obtain $\{\tilde{T}_{i}\}_{\{i\in\mathcal{I}_{2}\}}$ from Algorithm 1 applied with $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, $\mathcal{G}$, $\{C_{i}\}_{i\in{\mathcal{I}_{2}}}$.
6:return $\{(\tilde{T}_{i},C_{i})\}_{i\in\mathcal{I}_{2}}$.
Algorithm 3 Calibration with a known censoring mechanism
0: Calibration data $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, generative model $\mathcal{G}(\cdot)$, audit function $\texttt{Audit}(\cdot)$, pre-trained quantile regression model $\{\hat{q}_{\tau}(\cdot)\}_{\tau\in\mathcal{T}}$, target miscoverage rate $\alpha$}, prior quantile $\tau_{\textup{prior}}$}, calibration mode (Naive, Basic, Trimmed, Optimized), quantile trimming threshold $M$, total budget $B$.
1:if calibration mode is Naive then
2:  Sample $C_{i}\sim\mathrm{Geom}(n/B)$ for all $i\in{\mathcal{I}_{2}}$}
3:  Obtain $\{\tilde{T}_{i}\}_{\{i\in\mathcal{I}_{2}\}}$ from Algorithm 1 applied with $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$, $\mathcal{G}$, $\{C_{i}\}_{i\in{\mathcal{I}_{2}}}$.
4:  $w_{\tau}(X_{i})\leftarrow\frac{1}{\mathbb{P}[\hat{q}_{\tau}(X_{i})\leq C_{i}|X_{i}]},\ i\in{\mathcal{I}_{2}}$. // compute the weights for the geometric $C_{i}$.
5:else
6:  if calibration mode is Basic then
7:   $\{\hat{f}_{\tau}(X_{i})\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}\leftarrow\{\hat{q}_{\tau}(X_{i})\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}$
8:   $\{\pi_{i}\}\leftarrow\{\min(\tfrac{B}{|{\mathcal{I}_{2}}|\;\hat{f}_{\tau_{\textup{prior}}}(X_{i})},1)\}$ // per-prompt evaluation probability (Section 5.2)
9:  else if calibration mode is Trimmed then
10:   $\{\hat{f}_{\tau}(X_{i})\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}\leftarrow\{\min{(\hat{q}_{\tau}(X_{i}),M)}\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}$ // trim the quantile est. (Section 5.3)
11:   $\{\pi_{i}\}\leftarrow\{\min(\tfrac{B}{|{\mathcal{I}_{2}}|\;\hat{f}_{\tau_{\textup{prior}}}(X_{i})},1)\}$ // per-prompt evaluation probability (Section 5.2)
12:  else if calibration mode is Optimized then
13:   $\{\hat{f}_{\tau}(X_{i})\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}\leftarrow\{\min{(\hat{q}_{\tau}(X_{i}),M)}\}_{\tau\in\mathcal{T},i\in{\mathcal{I}_{2}}}$ // trim the quantile est. (Section 5.3)
14:   $\{\pi_{i}\}\leftarrow$ Algorithm 4 applied with $\{\hat{f}_{\tau_{\textup{prior}}}(X_{i})\}_{i\in{\mathcal{I}_{2}}}$ // optimized per-prompt evaluation                       probability (Section 5.4)
15:  end if
16:  $\{(\tilde{T}_{i},C_{i})\}_{i\in{\mathcal{I}_{2}}}\leftarrow$ Algorithm 2 applied to $\{X_{i}\}_{i\in{\mathcal{I}_{2}}}$ with $\{\pi_{i}\}_{i\in{\mathcal{I}_{2}}}$ and $\{\hat{f}_{\tau_{\textup{prior}}}(X_{i})\}_{i\in{\mathcal{I}_{2}}}$
17:  $w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\leftarrow\frac{1}{\pi_{i}},\ i\in{\mathcal{I}_{2}}$
18:end if
19:for $\tau\in\mathcal{T}\cap[0,{\tau_{\textup{prior}}}]$ do
20:  $\hat{\alpha}(\tau)\leftarrow\displaystyle\frac{1}{|{\mathcal{I}_{2}}|}\sum_{i\in\mathcal{I}_{2}}w(\{X_{j}\}_{j\in{\mathcal{I}_{2}}},i)\;\mathbb{I}\bigl\{\tilde{T}_{i}<\hat{f}_{\tau}(X_{i})\leq C_{i}\bigr\}$ // miscoverage est.
21:end for
[…]
23:return Lower predictive bound (LPB) for a test point $X_{\textup{test}}=x$, given by $\hat{L}(x)=\hat{f}_{\hat{\tau}}(x)$
@@
7:  $\lambda^{*}\leftarrow$ bisection on $\lambda_{\textup{low}},\lambda_{\textup{high}}]$ to solve $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\,\pi_{i}^{*}(\lambda)=B$ within $\epsilon$
@@ Appendix C A deeper look into $\hat{\alpha}(\tau)$
In Section 4, we introduced the miscoverage estimator $\hat{\alpha}(\tau)$. In this section, we explore its statistical properties in detail. We begin by deriving its expectation and variance:
@@
See Appendix D.1 for the proof. We remark that the conditional independence and pair-wise independence assumptions in Proposition C.1 hold for all variants of our method from Sections 5. This proposition shows that $\hat{\alpha}(\tau)$ is an unbiased miscoverage estimator, resulting in a useful approximation for the miscoverage estimation variance. Under a simplified assumption that, for a fixed $\tau$, each calibration point $i\in{\mathcal{I}_{2}}$ is miscovered at the same constant rate conditional on $\{X_{j}\}_{j\in\mathcal{I}_{2}}$, we have the following result.
@@ Appendix D Proofs
Observe that if $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\leq B$, then $\pi_{i}^{*}=1$ for all $i\in{\mathcal{I}_{2}}$ is the optimal solution which satisfies the constraint. We now turn to consider the case where $\sum_{i\in\mathcal{I}_{2}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$. We first show that there exists a unique solution $\lambda^{*}$. Observe that $U:(0,\infty)\to(0,\sum_{i}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})]$ is continuous and strictly decreasing, with $\lim_{\lambda\to 0^{+}}U(\lambda)=\sum_{i}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$ and $\lim_{\lambda\to\infty}U(\lambda)=0$. By the intermediate‐value theorem there is a unique $\lambda^{*}>0$ such that $U(\lambda^{*})=B$. Next, the output of Algorithm 4 is indeed the solution $\lambda^{*}$ since $U$ is a monotone function, and thus the bisection approach is valid in this case. Therefore, $\pi^{*}(\lambda)$ is the correct optimal solution. ∎
@@
Inductive step. We now suppose that the claim holds for any weight vector with a sum $B-1<S\leq M|{\mathcal{I}_{2}}|$, and show that it holds for any vector with a sum S1S-1. Given a weight vector $\{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\}_{i\in{\mathcal{I}_{2}}}$ with a sum $\sum_{i\in{\mathcal{I}_{2}}}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})=S-1$, we compose a new vector with a sum $S$, as follows. In $\{\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})\}_{i\in{\mathcal{I}_{2}}}$, there exists an index $i_{0}$ such that $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i_{0}})\leq M-1$. The new vector with a sum $S$ is defined by:
@@
Since $U^{\prime}$ is continuous and decreasing on $(0,\infty)$ with $\lim_{\mu\to 0^{+}}U^{\prime}(\mu)=\sum_{i}\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})>B$ and $\lim_{\mu\to\infty}U^{\prime}(\mu)=0$, the intermediate‐value theorem guarantees a unique $\lambda^{\prime}$ with $U^{\prime}(\lambda^{\prime})=B$. Moreover, since $U^{\prime}$ is decreasing, and $U^{\prime}(\lambda)\geq B=U^{\prime}(\lambda^{\prime})$ we get $\lambda^{\prime}\geq\lambda$. Define $\pi_{i}$ the optimal solution of (10) with $\hat{f}_{{\tau_{\textup{prior}}}}(X_{i})$ and $\pi^{\prime}_{i}$ the optimal solution with $\hat{f}^{\prime}_{{\tau_{\textup{prior}}}}(X_{i})$.